OpenStudy (anonymous):

A number was added to each value in an original data set to give a new data set with values 14 , 2,13 , 4, and 2. What number was added to each value if the original mean was 6

OpenStudy (anonymous):

14 + 2 + 13 + 4 + 2 = 35/6 is new mean ------------------ 6 original mean =6 then 6 +x = 35/6 x = 35 - 36 ------ 6 x = -1/6

OpenStudy (anonymous):

-1/6 is the number.

OpenStudy (anonymous):

here are the multiple choice ansewers 1 4 7 8

OpenStudy (anonymous):

but the number should be negative the new mean is 35/6 which is less than 36/6=6 check your numbers again maybe you must have given me wrong numbers due to some typo .

OpenStudy (anonymous):

ok um here ill re copy and paste it A number was added to each value in an original data set to give a new data set with values ,14 ,2,13 , 4, and 2. What number was added to each value if the original mean was 6